OpenStudy (anonymous):

:) Find the measures of two angles, one positive and one negative, that are coterminal with pi divided by two.

OpenStudy (anonymous):

Subtract and add it to \(\huge 2\pi\) or \(\huge\frac{ 4\pi }{ 2 }\).